OpenStudy (anonymous):

find the range of possible values of x for WYXV. VW is 5.05. XV is 5.05. WY is 10.4 and YX is 2x-7. I dont know where to begin :(

OpenStudy (danjs):

draw a picture

OpenStudy (anonymous):



OpenStudy (danjs):

Find the length of VY by using law of cosines maybe

OpenStudy (danjs):

10.4^2 = 5.05^2 + [VY]^2 -2*5.05*[VY]*cos(92) Or Law of Sines may be easier sin(92)/10.4 = sin(Y)/5.05 = sin(W) / VY Find angle Y, then angle W = 180 - Y - 92

OpenStudy (anonymous):

I dont understand after I get the angle. Do I put it into an equation 5.05 + 2x-7 =180?

OpenStudy (danjs):

Find the Length of VY, then use law of cosines to find x, (2x-7)^2 = 5.05^2 + [VY]^2 - 2(5.05)*(VY)*cos(57)

OpenStudy (danjs):

Is that the topic you are studying now, law of sines/cosines?

OpenStudy (anonymous):

Not at all- we are studying area of similar figures theorems :(